# SelectLLM: Can LLMs Select Important Instructions to Annotate?

### 29 Jan 2024 v1

@@ Abstract
Training large language models (LLMs) with a large and diverse instruction dataset aligns the models to comprehend and follow human instructions. Recent works have shown that using a small set of high-quality instructions can outperform using large yet more noisy ones. Because instructions are unlabeled and their responses are natural text, traditional active learning schemes with the model’s confidence cannot be directly applied to the selection of unlabeled instructions. In this work, we propose a novel method for instruction selection, called SelectLLM, that leverages LLMs for the selection of high-quality instructions. Our high-level idea is to use LLMs to estimate the usefulness and impactfulness of each instruction without the corresponding labels (i.e., responses), via prompting. SelectLLM involves two steps: dividing the unlabelled instructions using a clustering algorithm (e.g., CoreSet) to multiple clusters, and then prompting LLMs to choose high-quality instructions within each cluster. SelectLLM showed comparable or slightly better performance on the popular instruction benchmarks, compared to the recent state-of-the-art selection methods. All code and data are publicly available [^1] https://github.com/minnesotanlp/select-llm.
@@ 1 Introduction
Instruction-tuning, which fine-tunes language models (LLMs) to follow the instructions constructed from a massive number of diverse and different tasks, has shown impressive generalization performance on various unseen tasks (Wei et al., 2022; Chung et al., 2022). However, creating large and diverse annotated instruction datasets requires the huge cost of human labeling, and hence it becomes a major challenge for instruction-tuning. While synthetic datasets generated by advanced large language models (LLMs) like GPT-4 have partly addressed this issue (Taori et al., 2023; Wang et al., 2022a), they often contain low-quality data, highlighting the need for more focus on dataset refinement (Zhou et al., 2023; Cao et al., 2023).
@@
Figure 1: Conceptual comparison between previous approaches to select instructions and SelectLLM. Focusing on input instructions (top) is unable to consider the difficulty or uncertainty of response. Response-based approaches (middle) can suffer from the inference cost and quality issues of synthetic responses. SelectLLM (bottom) does not suffer from these issues by estimating the effectiveness of instructions via prompting LLMs. More discussions are in Appendix B.
@@
A typical solution to address these challenges is to allocate human resources, i.e., developing a way to select important instructions after being labeled, similar to active learning (Settles, 2009). Since it is relatively easy to access unlabelled instructions or texts from various sources, one can create a high-quality small dataset if unlabelled instructions are well-selected and annotated accordingly. Previous selection algorithms usually consider density at the embedding space (Sener and Savarese, 2018) or uncertainty (or difficulty) of the generated predictions by the model (Kung et al., 2023); however, these methods are limited by being unable to incorporate difficulty of each sample or the cost of generating the inferences and quality issues from using synthetic labels, respectively (Figure 1). On the other hand, a few recent works show the capability of LLMs for measuring the quality or relevance of multiple texts (Sun et al., 2023; Liu et al., 2023). Inspired by these, we aim to explore a novel way to select unlabelled instructions by leveraging LLMs without inferences of their possible responses.
@@
Contribution. In this work, we introduce SelectLLM, a method that selects an effective subset of unlabelled instructions by prompting LLMs. Our high-level idea is using LLMs to estimate the usefulness and impact of each instruction without the corresponding labels; through the initial experiments presented in Figure 3, we first verify that LLMs do possess such a capability. To enhance the effectiveness of instruction selection via LLMs, SelectLLM first divides the entire dataset into several small subsets to construct the input queries for prompting; specifically, we use equal-size K-means clustering to create each subset with diverse unlabelled instructions while preserving the global structure of the entire dataset. Then, SelectLLM constructs an input query for each subset using a carefully designed input prompt for selecting with LLMs. Finally, we forward these constructed queries into LLMs, and SelectLLM selects a few instructions that are expected to be the most helpful in fine-tuning LLMs.
@@
We demonstrate the effectiveness of SelectLLM compared to various state-of-the-art selection methods on two popular benchmarks for instruction-tuning, Dolly (Conover et al., 2023) and Cleaned Alpaca (Taori et al., 2023). In the experiments, we verified that our method consistently outperforms the baselines across the varied selection sizes. For instance, when selecting samples from Dolly, SelectLLM exhibited nearly 2.5% - 3% relative improvements compared to the strongest baseline on the Rouge F1 score and cosine similarity, respectively across all sample sizes (1k/2k/3k). In addition, SelectLLM shows better cross-task generalizations, an important characteristic of instruction-tuned LMs. More interestingly, we observe that the generated responses by SelectLLM are preferred to the responses from other baselines when evaluated with GPT-4.
While the input prompt is designed to focus on improving the generic performance of LLMs, SelectLLM has flexibility that could be easily tuned for the specific purpose of the users, e.g., reducing the toxicity. Hence, we believe that our framework could open up the potential of LLMs for selecting the samples with any desired property, which could be flexibly set depending on the specific use case.
@@ 2 Related Work
Instruction-tuning for LMs. Instruction-tuning (Wei et al., 2022), a form of fine-tuning LLMs, has emerged as a prominent methodology to align pre-trained LMs for various tasks by describing their descriptions into the common form of instruction. Due to its ease of implementation and remarkable generalization capabilities for unseen tasks (Wei et al., 2022; Chung et al., 2022; Jang et al., 2023), it has gained substantial popularity recently. Constructing these instruction datasets with human annotations is a standard way (Conover et al., 2023; Sanh et al., 2021; Wang et al., 2022b), but it has a certain difficulty in terms of variety of instructions and the total number of instances due to labeling cost (Wang et al., 2022a); One promising solution for this limitation is to synthesize existing datasets and create diverse, multi-task datasets with the help of LLMs such as Alpaca (Taori et al., 2023) and Self-instruct (Wang et al., 2022a). However, using LLM-creating data increases the risk of including low-quality examples, and it is known that removing such noise from the dataset is critical for the effective instruction-tuning of LLMs (Zhou et al., 2023; Cao et al., 2023); therefore, in this work, we explore an alternative way to use LLMs for constructing a high-quality instruction dataset, by using LLMs for selecting unlabeled instructions.
@@
Sample selection for efficient instruction-tuning. Expanding instruction dataset comes with its own set of challenges, including resource and time constraints, the need for human annotation, and the prevalence of redundant data. Incorporating human efforts is one of the most straightforward ways to address these limitations; for example, Zhou et al. (2023) manually annotate and filter out the noisy ones using human annotators to select high-quality instructions. However, it could require huge costs when all the procedures are done by human labor; collecting abundant unlabelled instructions, and then selecting and annotating the useful ones could be an effective way to mitigate these issues, similar to the popular task of active learning (Settles, 2009; Sener and Savarese, 2018). Here, the most important thing is selection criteria.
@@
One line of the works focuses on density to include diverse instructions; for example, Chen et al. (2023a) maps the unlabelled instructions into embedding space and conducts K-means clustering (Hartigan and Wong, 1979) and K-center greedy algorithm (Sener and Savarese, 2018) for the selection. Another line of the works focuses on uncertainty (or difficulty) of instructions measured with LLMs’ outputs; Kung et al. (2023) propose to measure the uncertainty by inducing the perturbation on the input instruction and see how the responses are changed. Unlike these approaches, we directly prompt LLMs with unlabelled instructions, to select the few examples expected to be useful in training LLMs; from LLMs’ capability of reasoning and generating useful responses, we assume that they could infer the impactfulness of the unlabelled instructions. On the other hand, one recent work (Chen et al., 2023b) has used the prompting of LLMs to construct a high-quality instruction dataset. However, our work is totally different as we focus on selecting the unlabelled instructions while (Chen et al., 2023b) focus on filtering low-quality labeled instructions.
@@ 3.1 Preliminary
We first describe the problem setup of our interest. Let $\mathcal{X}=\{x_{i}\}_{i=1}^{M}$ denote the given un-annotated dataset, where $x_{i}$ represents $i$th instruction example. Then, our goal is to select $N$ most effective instructions from $\mathcal{X}$ which will be labeled by human annotators, to fine-tune a given large language model (LLM) $f_{\theta}$, e.g., LLaMA-2 (Touvron et al., 2023), and make it be generalized for various instructions. Specifically, let’s assume that we select $N$ instructions under a criteria $s(j)$:
@@
where $S$ denotes the indices of selected instructions. Then, the selected instruction $x\in\mathcal{X}(S)$ is labeled with a corresponding response $y$ by human annotators, and it results in the annotated instruction dataset $\mathcal{D}$, where $\mathcal{D}=\{(x_{s(j)},y_{s(j)})\}_{j=1}^{N}$. 222 We assume that the human responses could be obtained right after the selection for the experiments. Therefore, the performance of fine-tuned LLMs on $\mathcal{X}(S)$ is significantly varied depending on which selection criteria $s(j)$ is considered. While various criteria have been explored under tasks like active learning (see Section 2), this direction is less explored under the regime of instruction-tuning.
@@ 3.2 Selection via prompting LLMs
For the selection criteria $s(j)$, SelectLLM proposes to leverage LLMs with a properly designed prompt without using ground truth or generated responses. Our high-level intuition is that LLMs can infer the potential impact of each instruction by only reading the instruction; as shown in Figure 3, we observed that the recent LLMs, e.g., ChatGPT, could estimate the effectiveness of each instruction for model training (e.g., LLaMA-2), even without the corresponding responses. To further improve the effectiveness of selection via LLMs, we carefully designed the input prompt to incorporate several important perspectives for instruction-tuning, and it is presented in 4. Formally, this process could be described as follow: we first assume that the dataset $\mathcal{X}$ is divided into $K$ smaller and non-overlapped subsets, i.e., $\mathcal{X}=\bigcup_{k=1}^{K}\mathcal{X}_{k}$. Then, we construct input query $q_{t}$ using the designed prompt $p_{\tt sel}$ and $\mathcal{X}_{k}$, and forward it to LLM to select $O=\lfloor N/K\rfloor$ examples:
$S_{t}=\texttt{LLM}\big{(}p_{\tt sel}(q_{t},O)\big{)}$ (2)
where $S_{t}=\{s_{t}(j)\}_{j=1}^{O},~{}s_{t}(j)\in[1,N]$.
@@ 3.3 Composing query of LLMs via clustering
In addition, to further improve the effectiveness of using LLMs for selection, we carefully design how to divide the entire dataset into several subsets which would be used to construct input queries, based on the equal-size clustering method. Here, our high-level idea is composing the subsets that maximize the diversity among the instructions while maintaining the global structure of the dataset as much as possible. Specifically, we first extract the sentence embedding of the instructions in $\mathcal{X}$, using the pre-trained sentence encoder $g_{\phi}$, e.g., Sentence-BERT (Reimers and Gurevych, 2019). Then, we conduct K-means clustering (Hartigan and Wong, 1979) on these embeddings, and calculate $D\in\mathbb{R}^{N\times K}$, the distance of all instances in $\mathcal{X}$ to $K$ cluster centers $c_{1},\cdots,c_{K}$. Based on the distances, we assigned each instance $x$ among $[1,K]$, by iteratively taking the one with the shorted distance to the cluster center among the remaining instances, to guarantee equal sizes for each $k$.
Overall, the selection procedure of SelectLLM is as follows: (1) constructing the input queries by separating the entire dataset into multiple subsets of diverse instructions. Then, (2) feed these queries into LLM and get the selected indices. The formal presentation of these procedures is presented in Algorithms 1 and 2 in Appendix.
@@ 4 Experiments
Table 1: Experimental results on Dolly (Conover et al., 2023). Rouge (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 by different numbers of examples are compared. The best and second best scores are highlighted in bold and underline, respectively.
Rouge (F1) Cosine Similarity Avg Across Sizes Methods $1k$ $2k$ $3k$ $1k$ $2k$ $3k$ Avg Rouge Avg Cosine Lengthshort 0.073 0.109 0.130 0.192 0.265 0.336 0.104 0.264 Perplexity 0.158 0.183 0.192 0.402 0.433 0.453 0.178 0.429 CBSsbert 0.147 0.200 0.216 0.359 0.473 0.512 0.188 0.448 Lengthlong 0.256 0.247 0.238 0.641 0.626 0.611 0.247 0.626 CBSinstr 0.258 0.255 0.255 0.617 0.638 0.632 0.256 0.629 Random 0.239 0.264 0.278 0.589 0.644 0.650 0.260 0.628 Diversity 0.237 0.275 0.282 0.582 0.650 0.666 0.265 0.633 OpenEnd 0.258 0.271 0.282 0.627 0.641 0.669 0.270 0.646 Coreset 0.271 0.281 0.279 0.649 0.662 0.659 0.277 0.657 Ours 0.278 0.288 0.289 0.668 0.680 0.686 0.285 0.678
@@
    Does SelectLLM outperform the previous state-of-the-art selection methods for instruction-tuning LLMs? (Tables 1,  2, and 3)
@@
    Do LLMs tuned with SelectLLM provide good responses qualitatively? (Tables 4 and 6)
@@
  • $\circ$
    Why SelectLLM is more effective than other selection methods ? (Table 7)

### 20 Feb 2024 v2

@@ Abstract
Instruction tuning benefits from large and diverse datasets, however creating such datasets involves a high cost of human labeling. While synthetic datasets generated by large language models (LLMs) have partly solved this issue, they often contain low-quality data. One effective solution is selectively annotating unlabelled instructions, especially given the relative ease of acquiring unlabeled instructions or texts from various sources. However, how to select unlabelled instructions is not well-explored, especially in the context of LLMs. Further, traditional data selection methods, relying on input embedding space density, tend to underestimate instruction sample complexity, whereas those based on model prediction uncertainty often struggle with synthetic label quality. Therefore, we introduce SelectLLM, an alternative framework that leverages the capabilities of LLMs to more effectively select unlabeled instructions. SelectLLM consists of two key steps: Coreset-based clustering of unlabelled instructions for diversity and then prompting a LLM to identify the most beneficial instructions within each cluster. Our experiments demonstrate that SelectLLM matches or outperforms other state-of-the-art methods in instruction tuning benchmarks. It exhibits remarkable consistency across human and synthetic datasets, along with better cross-dataset generalization, as evidenced by a 10% performance improvement on the Cleaned Alpaca test set when trained on Dolly data. All code and data are publicly available [^1] https://github.com/minnesotanlp/select-llm.
@@ 1 Introduction
Instruction tuning, which fine-tunes language models (LMs) to follow instructions constructed from a massive number of diverse and different tasks, has shown impressive generalization performance on various unseen tasks (Wei et al., 2022; Chung et al., 2022). However, creating large and diverse annotated instruction datasets is a major challenge due to the significant cost of human labeling. While synthetic datasets generated by advanced large language models (LLMs) have partly addressed this issue (Taori et al., 2023; Wang et al., 2022a), they often contain low-quality data, highlighting the need for more focus on dataset refinement (Zhou et al., 2023; Cao et al., 2023).
@@
Figure 1: Conceptual comparison between previous approaches to select instructions and SelectLLM. Focusing on input instructions (top) is unable to consider the difficulty or uncertainty of response. Output-based methods (middle) can suffer from the inference cost and quality issues of synthetic responses. SelectLLM (bottom) does not suffer from these issues by estimating the effectiveness of instructions via prompting LLMs.
@@
As it is relatively easy to access unlabelled instructions or texts from various sources, one promising solution to address these challenges is developing a way to select important unlabelled instructions for annotation, similar to active learning (Settles, 2009); from this, one can create a high-quality small dataset if unlabelled instructions are well-selected and annotated accordingly. However, existing selection methods present limitations. Active learning-based algorithms focusing on density in the embedding space (Sener and Savarese, 2018) often overlook the complexity of instruction samples. Alternatively, output-based approaches that assess uncertainty in model predictions (Kung et al., 2023) grapple with high computational costs and the challenges posed by synthetic label quality (Figure 1). In light of these challenges, a few recent works show the capability of LLMs for measuring the quality or relevance of multiple texts (Sun et al., 2023; Liu et al., 2023). Inspired by these, we investigate the potential of LLMs to select effective unlabelled instructions, leveraging their vast knowledge base to discern the complexity and utility of each instruction without generating inferences.
@@
Contribution. In this work, we introduce SelectLLM, a method that selects an effective subset of unlabelled instructions by prompting LLMs. At a high level, our method uses LLMs to estimate the usefulness and impact of each instruction without the corresponding labels; through the initial experiments presented in Figure 3, we first verify that LLMs do possess such a capability. To improve instruction selection via LLMs, SelectLLM first divides the entire dataset into several small subsets to construct the input queries for prompting; specifically, we use equal-size K-means clustering to create each subset with diverse unlabelled instructions while preserving the overall dataset structure. Then, SelectLLM constructs an input query for each subset using a carefully designed input prompt for selecting with LLMs. Finally, we forward these constructed queries into the LLM, and SelectLLM selects a few instructions that are expected to be the most helpful in fine-tuning models.
@@
We demonstrate the effectiveness of SelectLLM compared to various state-of-the-art selection methods on two popular benchmarks for instruction tuning, Dolly (Conover et al., 2023) and Cleaned Alpaca (Taori et al., 2023). Our experiments show that SelectLLM consistently outperforms the baselines across the varied selection sizes. For instance, when selecting samples from Dolly, SelectLLM exhibited nearly 2.5% - 3% relative improvements compared to the strongest baseline on the Rouge-L F1 score and cosine similarity, respectively across all sample sizes (1k/2k/3k). In addition, SelectLLM shows better cross-task generalizations, an important characteristic of instruction-tuned LMs. More interestingly, we observe that the generated responses by SelectLLM are preferred to the responses from other baselines when evaluated with GPT-4.
Additionally, the input prompt in SelectLLM is primarily designed to select instructions that enhance the overall performance of LLMs. However, its flexible nature allows for easy customization to meet specific user needs (e.g., reducing toxicity), by simply adjusting SelectLLM’s input prompt during sample selection. This adaptability of our framework paves the way for using LLMs to select samples with various desired characteristics, tailored to suit distinct use cases, which is not possible with previous selection methods.
@@ 2 Related Work
Instruction tuning for LMs. Instruction tuning (Wei et al., 2022), a form of fine-tuning LLMs, has emerged as a prominent methodology to align pre-trained LMs for various tasks by describing the tasks in the common form of instructions. Due to its ease of implementation and remarkable generalization capabilities for unseen tasks (Wei et al., 2022; Chung et al., 2022; Jang et al., 2023), it has gained substantial popularity recently. Constructing these instruction datasets with human annotations is a standard way (Conover et al., 2023; Sanh et al., 2021; Wang et al., 2022b), but this method faces challenges in terms of variety of instructions and the total number of instances due to labeling cost (Wang et al., 2022a); one promising solution for this limitation is to synthesize existing datasets and create diverse, multi-task datasets with the help of LLMs such as Alpaca (Taori et al., 2023) and Self-instruct (Wang et al., 2022a). However, using LLM-created data increases the risk of including low-quality examples, and it is known that removing such noise from the dataset is critical for the effective instruction tuning of LLMs (Zhou et al., 2023; Cao et al., 2023); therefore, in this work, we explore an alternative way to use LLMs to construct a high-quality instruction dataset, by using LLMs to select unlabeled instructions.
@@
Sample selection for efficient instruction tuning. Expanding instruction datasets has its own set of challenges, including the need for extensive resources, time, human annotation, and the prevalence of redundant data. A common solution involves human intervention, such as the approach by Zhou et al. (2023), which manually annotates high-quality instructions after filtering out low-quality data. However, there are huge costs associated with such procedures done by human labor. An effective solution is to gather a large pool of unlabeled instructions and then selectively annotate the most useful ones, similar to the popular task of active learning (Settles, 2009; Sener and Savarese, 2018). Here, the most important thing is selection criteria.
@@
One line of work focuses on density to include diverse instructions; for example, Chen et al. (2023a) map the unlabelled instructions into an embedding space and use K-means clustering (Hartigan and Wong, 1979) and the K-center greedy algorithm (Sener and Savarese, 2018) for the selection. Another line of work focuses on uncertainty (or difficulty) of instructions measured with LLMs’ outputs; Kung et al. (2023) measure uncertainty by observing how LLM-generated responses vary with changes in the input instructions. Unlike these approaches, we directly prompt LLMs with unlabelled instructions, to select the few examples expected to help train the model; from LLM’s capability of reasoning and generating useful responses, we assume that they could infer the impactfulness of the unlabelled instructions. Meanwhile, a recent study by (Chen et al., 2023b) also prompts LLMs to construct a high-quality instruction dataset. However, our work differs as we focus on selecting unlabelled instructions while (Chen et al., 2023b) focuses on filtering out low-quality labeled instructions.
@@ 3.1 Preliminary
We first describe the problem setup of our interest. Let $\mathcal{X}=\{x_{i}\}_{i=1}^{M}$ denote the given unlabeled dataset, where $x_{i}$ represents $i$th unlabeled instruction and $M$ is the total number of instructions. Then, our goal is to select $N$ most effective instructions from $\mathcal{X}$ which will be labeled by human annotators, to fine-tune a target large language model (LLM) $f_{\theta}$, e.g., LLaMA-2 (Touvron et al., 2023), and make it be generalized for various instructions. Formally, we select $N$ instructions from $\mathcal{X}$ under a selection criteria $s(j)$:
@@
where $S$ denotes the indices of selected instructions. Then, the selected instruction $x\in\mathcal{X}(S)$ is labeled with a corresponding label $y$ by human annotators, and it results in the annotated instruction dataset $\mathcal{D}$, where $\mathcal{D}=\{(x_{s(j)},y_{s(j)})\}_{j=1}^{N}$. Therefore, the performance of LLM $f_{\theta}$ fine-tuned on $\mathcal{X}(S)$ significantly varies depending on which selection criteria $s(j)$ is considered. While various selection criteria have been explored under tasks like active learning (see Section 2), this direction is less explored under the paradigm of instruction tuning.
@@ 3.2 Selection via prompting LLMs
For the selection criteria $s(j)$, SelectLLM proposes to use LLMs with a properly designed prompt without using ground truths or generated labels. Our high-level intuition is that LLMs can infer the potential impact of each instruction by only reading the instruction; as shown in Figure 3, we observed that the recent LLM, e.g., ChatGPT, could estimate the effectiveness of each instruction for model training (e.g., LLaMA-2), even without the corresponding labels. To further improve the effectiveness of selection via LLMs, we carefully designed the input prompt to incorporate several important perspectives for instruction tuning, and it is presented in Figure 4. Formally, this process could be described as follows: we first assume that the dataset $\mathcal{X}$ is divided into $K$ non-overlapped subsets, i.e., $\mathcal{X}=\bigcup_{k=1}^{K}\mathcal{X}_{k}$. Then, we construct input query $q_{k}$ using the designed prompt $p_{\tt sel}$ and $\mathcal{X}_{k}$, and forward it to LLM to select $\widetilde{N}=\lfloor N/K\rfloor$ examples:
$S_{k}=\texttt{LLM}\big{(}p_{\tt sel}(q_{k},\widetilde{N})\big{)}$ (2)
where $S_{k}=\{s_{k}(j)\}_{j=1}^{\widetilde{N}},~{}s_{k}(j)\in[1,N]$.
@@ 3.3 Composing query of LLMs via clustering
To further improve the effectiveness of using LLMs for selection, we carefully design how to divide the entire dataset into several subsets which would be used to construct input queries, based on the equal-size clustering method. Here, our high-level idea is composing the subsets that maximize the diversity among the instructions while maintaining the global structure of the dataset. Specifically, we first extract the embeddings of the instructions in $\mathcal{X}$, using the pre-trained sentence encoder $g_{\phi}$ such as Sentence-BERT (Reimers and Gurevych, 2019a). Then, we conduct K-means clustering (Hartigan and Wong, 1979) on these embeddings, and calculate $D\in\mathbb{R}^{N\times K}$, the distance of all instances in $\mathcal{X}$ to $K$ cluster centers $c_{1},\cdots,c_{K}$. Based on the distances, we assigned each instance $x$ among $[1,K]$, by iteratively taking the one with the shorted distance to the cluster center among the remaining instances, to guarantee equal sizes for each $k$.
Overall, the selection procedure of SelectLLM is as follows: (1) construct input queries by separating the entire dataset into multiple subsets of diverse instructions. Then, (2) feed these queries into a LLM and get the selected indices. The formal presentation of these procedures is presented in Algorithms 1 and 2 in Appendix.
@@ 4 Experiments
Table 1: Experimental results on Dolly (Conover et al., 2023). Rouge-L (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 by different numbers of examples are compared. The best and second best scores are highlighted in bold and underline, respectively.
Rouge-L (F1) Cosine Similarity Avg Across Sizes Methods $1k$ $2k$ $3k$ $1k$ $2k$ $3k$ Avg Rouge Avg Cosine Lengthshort 0.073 0.109 0.130 0.192 0.265 0.336 0.104 0.264 Perplexity 0.158 0.183 0.192 0.402 0.433 0.453 0.178 0.429 CBSsbert 0.147 0.200 0.216 0.359 0.473 0.512 0.188 0.448 Lengthlong 0.256 0.247 0.238 0.641 0.626 0.611 0.247 0.626 CBSinstr 0.258 0.255 0.255 0.617 0.638 0.632 0.256 0.629 Random 0.239 0.264 0.278 0.589 0.644 0.650 0.260 0.628 Diversity 0.237 0.275 0.282 0.582 0.650 0.666 0.265 0.633 OpenEnd 0.258 0.271 0.282 0.627 0.641 0.669 0.270 0.646 Coreset 0.271 0.281 0.279 0.649 0.662 0.659 0.277 0.657 Ours 0.278 0.288 0.289 0.668 0.680 0.686 0.285 0.678
@@
    Does SelectLLM outperform the previous state-of-the-art selection methods for instruction tuning LLMs? (Tables 123) If so, why? (Table 7)
@@
    Do LLMs tuned with SelectLLM provide good responses qualitatively? (Tables 46)